# Spatial-temporal Memories Enhanced Graph Autoencoder for Anomaly Detection in Dynamic Graphs

### v1

@@ V-B Baselines
To validate the effectiveness of STRIPE, we conducted a comparative analysis with six state-of-the-art node anomaly detection baselines. This comparison includes two dynamic node anomaly detection methods: NetWalk [6] and MTHL [12]. Given the limited number of dynamic node anomaly detection baselines, we also incorporate four of the most advanced static node anomaly detection methods in our comparison: DOMINANT [36], CoLA [11], SL-GAD [37] and GRADATE [38]. Details of these baselines are introduced as follows:
@@
TABLE III: Node anomaly detection performance on six benchmark datasets, the best and second to best results on each dataset are in bold and underlined, respectively. PRE, F1, and AUC represent the precision, macro-F1, and area under the curve, respectively.
Dataset DBLP-3 DBLP-5 Reddit Brain
Metrics PRE F1 AUC PRE F1 AUC PRE F1 AUC PRE F1 AUC
DOMINANT 0.1358 0.5490 0.6091 0.5431 0.7327 0.9154 0.4912 0.6326 0.8717 0.5845 0.5958 0.8212
CoLA 0.4753 0.4874 0.5814 0.4750 0.4872 0.4806 0.4747 0.4870 0.2410 0.4760 0.4877 0.5716
SL-GAD 0.5302 0.5193 0.6174 0.5229 0.5038 0.7638 0.4908 0.4844 0.6843 0.6640 0.6347 0.8735
GRADATE 0.5126 0.4874 0.5581 0.4923 0.4875 0.4726 0.5231 0.4870 0.5823 0.4910 0.4871 0.5629
NetWalk 0.5336 0.5092 0.7126 0.5805 0.5034 0.9107 0.5606 0.5726 0.7821 0.6239 0.6643 0.8590
MTHL 0.5758 0.5077 0.5901 0.5433 0.4951 0.7382 0.6295 0.6279 0.7074 0.5843 0.5988 0.8119
(a) DBLP-3
(b) DBLP-5
(c) Reddit
(d) Brain
Figure 4: ROC curves on four benchmark datasets for node anomaly detection.
V-C Metrics
To assess the performance of STRIPE and its competitors, we utilize a combination of ROC-AUC, precision, and Macro-F1 as evaluation metrics. ROC-AUC, a standard metric in anomaly detection, involves plotting the true positive rate against the false positive rate, with AUC (Area Under the Curve) representing the area under this ROC curve, ranging between 0 and 1. Higher AUC values indicate better performance. Additionally, we use precision and macro-F1 as our evaluation metrics. Precision measures the proportion of positive identifications that were actually correct. Macro-F1, as the harmonic mean of precision and recall, is computed individually for each class and then averaged, to provide a balanced measure of the model’s performance.
V-D Parameter Settings
@@
All the parameters can be tuned by 5-fold cross-validation on a rolling basis. We set the time window size $\tau$ as 3 and temporal convolution kernel width $K_{t}$ as 2. Both the GCN and gated temporal convolution encoders/decoders have 2 layers with hidden dimensions set as 32 for Brain and 128 for the remaining datasets. Balance factor $\alpha$ is set as 0.9 for Brain and Reddit, and 0.3 for DBLP-3 and DBLP-5. The training epoch is 20 and the learning rate is 0.001 for all the datasets. Evaluation round $R$ is 40 for Brain and 20 for the other datasets. The number of spatial and temporal memory items $P_{s}$ and $P_{t}$ are both set as 6. For the evaluation of static graph anomaly detection baselines, we first train these methods and measure the anomaly scores for each graph snapshot, then we derive the final score for evaluation by averaging these anomaly scores across all snapshots.
TABLE IV: Quantitative results w.r.t. precision, recall and AUC for ablation study
@@
V-E Comparison with the State-of-the-art Baselines
To evaluate the effectiveness of STRIPE on dynamic node anomaly detection, we compare its performance with six state-of-the-art baselines on four benchmark datasets. The precision, macro-F1 and AUC values are presented in Table III. The ROC curves can be found in Fig. 4. All the differences between our model and others are statistically significant ($p<0.01$). According to these results, we have the following findings:
[…]
V-F Ablation Study
To further investigate the contribution of each component in STRIPE, we perform an ablation study in this section. The results are shown in Table IV. We set five variants of STRIPE: w/o attribute, w/o structure, w/o temporary, w/o prototype and STRIPE. Among them, w/o attribute excludes the attributive reconstruction error by setting $\alpha$=0. w/o structure excludes the structural reconstruction error by setting $\alpha$=1. w/o temporary excludes the gated temporal convolution component and directly applies static version of STRIPE on each graph snapshot and computes the average anomaly scores across all the snapshots. w/o s-prototype and w/o t-prototype exclude the spatial and temporal memory modules, respectively, and reconstructs graphs without fusing prototypical patterns.
@@ V-F Ablation Study
(b) DBLP-5
(c) Reddit
(d) Brain
Figure 6: The sensitivity of spatial memory item number $P_{s}$ and temporal memory item number $P_{t}$ on four datasets. The vertical axis represents the AUC values of STRIPE with different $P_{s}$ and $P_{t}$. A darker color indicates a higher AUC value.
The comparison between w/o attribute, w/o structure, and STRIPE, as illustrated in Table IV, reveals a decline in performance when either attributive or structural reconstruction error is excluded from the loss function. This indicates that incorporating both types of reconstruction is essential for optimal performance in the node anomaly detection task. Furthermore, removing the gated temporal convolution component, as shown in the w/o temporary comparison with STRIPE, results in the most pronounced decrease in performance. This underscores the significance of tracking graph evolution for dynamic graph anomaly detection. Additionally, the performance difference between w/o s-prototype and STRIPE and the performance gap between w/o t-prototype and STRIPE further illustrate the contribution of both spatial and temporal normality memory items on the enhancement of anomaly detection performance.
@@
(b) DBLP-5
(c) Reddit
(d) Brain
Figure 7: The sensitivity of temporal convolution size $K_{t}$ and time window size $\tau$ on four datasets. The vertical axis represents the AUC values of STRIPE with different $K_{t}$ and $\tau$. A darker color indicates a higher AUC value.
@@
V-G Parameter Sensitivity
V-G1 Training ratio
V-G2 Number of memory items
The performance of our model, STRIPE, exhibits variability with adjustments in $P_{s}$ and $P_{t}$ across the DBLP-3, DBLP-5, and Brain datasets. Conversely, the Reddit dataset shows minimal performance fluctuation in response to changes in these parameters, likely due to its simpler normality patterns. This suggests that STRIPE can effectively capture normal spatial and temporal patterns with a minimal number of memory items, maintaining robust anomaly detection performance in simpler datasets.
For the DBLP-3, DBLP-5, and Brain datasets, we observe that increasing $P_{s}$ or $P_{t}$ from lower values generally leads to higher AUC scores. For example, enhancing $P_{t}$ from 2 to 6 with $P_{s}$ fixed at 2 in DBLP-3 improves performance, underscoring that a sparse memory set may not adequately represent the graph’s complex patterns. Nonetheless, further increasing the count of memory items beyond a certain point does not always yield better performance; it may in fact impair detection capabilities. For instance, elevating $P_{s}$ from 6 to 10 while keeping $P_{t}$ at 2 in Brain illustrates this trend, suggesting that excessively large memory modules might overemphasize specific details over general normality patterns, enabling abnormality to be reconstructed accurately.
V-G3 Parameters of temporal convolution
V-G4 Reconstruction Weight
V-G5 Hidden Dimension
In this experiment, we explored how varying the hidden dimension $D^{\prime}$ of both the encoder and decoder affects performance by adjusting $D^{\prime}$ from 8 to 256 and observing the AUC values. As shown in 8(b), we found that increasing $D^{\prime}$ enhances model performance within the range of [8, 128] for the DBLP-3, DBLP-5, and Reddit datasets, and within [8, 32] for Brain. Beyond these ranges, performance gains were negligible or even negative. Consequently, we opted for a $D^{\prime}$ of 32 for Brain and 128 for the other datasets.
V-G6 Evaluation Rounds
@@
V-H Case Study
In this section, we demonstrate the effectiveness of our spatial and temporal memory modules by visualizing the compactness loss between memory items with both normal and abnormal nodes. We configured the spatial and temporal memory items to six each and chose three normal and three abnormal nodes from the DBLP-3 test set for analysis. The compactness loss is calculated as outlined in Eq. 17. The findings are presented in Fig. 9, with darker colors indicating higher compactness loss and vice verca.
From Figs. 9(a) and 9(b), we note two key observations: (1) Normal nodes exhibit proximity to only a subset of memory items, showing distance from the rest. For instance, Fig. 9(a) illustrates that the compactness loss between the first normal node and the third spatial memory item is merely 0.06, considerably less than its loss with other memory items. This underscores the memory items’ capacity to distinctively and effectively encapsulate normal patterns, even with a limited set of items, while ensuring sufficient separation among them. (2) Conversely, abnormal nodes display substantial compactness losses across all memory items in both spatial and temporal dimensions. As an example, the compactness losses between the second abnormal node and all memory items exceed 0.89, highlighting the anomaly nodes’ divergence from established normal patterns. This deviation contributes to increased reconstruction errors and, consequently, elevated anomaly scores, effectively signaling their anomalous nature.

### v2

@@ V-B Baselines
To validate the effectiveness of STRIPE, we conducted a comparative analysis with nine state-of-the-art node anomaly detection baselines. This comparison includes five dynamic node anomaly detection methods: NetWalk [6], MTHL [12], TADDY [34], SAD [15] and CLDG [16]. Given the limited number of dynamic node anomaly detection baselines, we also incorporate four of the most advanced static node anomaly detection methods in our comparison: DOMINANT [22], CoLA [11], SL-GAD [24] and GRADATE [48]. Details of these baselines are introduced as follows:
@@
  • TADDY [34] develops coupled spatial-temporal patterns via a dynamic graph transformer to detect anomalies in graph streams.
  • SAD [15] is a semi-supervised dynamic graph anomaly detection model that combines memory networks with pseudo-label contrastive learning.
  • CLDG [16] is a contrastive learning model that performs representation learning on both discrete-time and continuous-time dynamic graphs.
V-C Parameter Settings
@@
All the parameters can be tuned by 5-fold cross-validation on a rolling basis. We set the time window size $\tau$ as 3 and temporal convolution kernel width $K_{t}$ as 2. Both the GCN and gated temporal convolution encoders/decoders have 2 layers with hidden dimensions set as 32 for Brain and 128 for the remaining datasets. Balance factor $\alpha$ is set as 0.9 for Brain and Reddit, and 0.3 for DBLP-3, DBLP-5 and OTC. The training epoch is 20 and the learning rate is 0.001 for all the datasets. Evaluation round $R$ is 40 for Brain and 20 for the other datasets. The number of spatial and temporal memory items $P_{s}$ and $P_{t}$ are both set as 6. For the evaluation of static graph anomaly detection baselines, we first train these methods and measure the anomaly scores for each graph snapshot, then we derive the final score for evaluation by averaging these anomaly scores across all snapshots.
V-D Comparison with the State-of-the-art Baselines
To evaluate the effectiveness of STRIPE on dynamic node anomaly detection, we compare its performance with six state-of-the-art baselines on four benchmark datasets. The precision, macro-F1 and AUC values are presented in Table II. All the differences between our model and others are statistically significant ($p<0.01$). According to these results, we have the following findings:
[…]
Figure 4: Anomaly detection under different anomaly rates on DBLP-3 (left) and Reddit (right) datasets.
V-E Performance under Different Anomaly Rates
To further evaluate the robustness of the proposed model, we further compare the performance of STRIPE with other baselines under different anomaly rates. The anomaly rate is the proportion of anomalous nodes to the total number of nodes, which can be changed by varying the number of injected anomalies $q$ to synthetic datasets. We vary anomaly rate from 0.01 to 0.05 and the result is reported in Fig. 4. We can observe that (1) STRIPE consistently outperforms other baselines across all the anomaly rates, which demonstrate the robustness and generalization of our model. (2) STRIPE generally has larger performance gain over other baselines under higher anomaly rates, and only has marginal performance drop compared to baselines such as CLDG or SL-GAD.
Figure 5: Evaluation of time efficiency. Left: The linear increase of both training and inference time of STRIPE w.r.t. node numbers in DGraph dataset. Right: Comparison of training and inference time of STRIPE with three most competitive baselines on DGraph dataset.
V-F Model Efficiency Evaluation
In this subsection, we provide experimental analysis of the computational efficiency of the proposed STRIPE. We measure the running time per epoch of STRIPE on the DGraph dataset while varying the number of nodes. As depicted in the left subplot of Fig. 5, both the training and inference time is linear w.r.t. node numbers, which aligns with our analysis in section IV-E. Additionally, we compare the time efficiency of STRIPE with baselines that acquire AUC scores over 60% on DGraph dataset, as shown in the right subplot of Fig. 5. Compared to CLDG, the second most accurate model following STRIPE, despite STRIPE is 2.86$\times$ slower than CLDG during inference, STRIPE is 4.62$\times$ faster in training and demonstrates a significant improvement of 5.8% in AUC scores compared to CLDG. Furthermore, STRIPE is 23.76$\times$ faster in training and 56.8$\times$ faster in inference compared to the second most accurate baseline, SAD.
TABLE III: Quantitative results w.r.t. precision, recall and AUC for ablation study
@@ V-F Model Efficiency Evaluation
V-G Ablation Study
To further investigate the contribution of each component in STRIPE, we perform an ablation study in this section. The results are shown in Table III. We set five variants of STRIPE: w/o attribute, w/o structure, w/o temporary, w/o s-prototype, w/o t-prototype and STRIPE. Among them, w/o attribute excludes the attributive reconstruction error by setting $\alpha$=0. w/o structure excludes the structural reconstruction error by setting $\alpha$=1. w/o temporary excludes the gated temporal convolution component and directly applies static version of STRIPE on each graph snapshot and computes the average anomaly scores across all the snapshots. w/o s-prototype and w/o t-prototype exclude the spatial and temporal memory modules, respectively, and reconstructs graphs without fusing prototypical patterns.
The comparison between w/o attribute, w/o structure, and STRIPE, as illustrated in Table III, reveals a decline in performance when either attributive or structural reconstruction error is excluded from the loss function. This indicates that incorporating both types of reconstruction is essential for optimal performance in the node anomaly detection task. Furthermore, removing the gated temporal convolution component, as shown in the w/o temporary comparison with STRIPE, results in the most pronounced decrease in performance. This underscores the significance of tracking graph evolution for dynamic graph anomaly detection. Additionally, the performance difference between w/o s-prototype and STRIPE and the performance gap between w/o t-prototype and STRIPE further illustrate the contribution of both spatial and temporal normality memory items on the enhancement of anomaly detection performance.
V-H Parameter Sensitivity
V-H1 Training ratio
V-H2 Number of memory items
The performance of our model, STRIPE, exhibits variability with adjustments in $P_{s}$ and $P_{t}$ across the DBLP-3 and Brain datasets. Conversely, the Reddit dataset shows minimal performance fluctuation in response to changes in these parameters, likely due to its simpler normality patterns. This suggests that STRIPE can effectively capture normal spatial and temporal patterns with a minimal number of memory items, maintaining robust anomaly detection performance in simpler datasets.
For the DBLP-3 and Brain datasets, we observe that increasing $P_{s}$ or $P_{t}$ from lower values generally leads to higher AUC scores. For example, enhancing $P_{t}$ from 2 to 6 with $P_{s}$ fixed at 2 in DBLP-3 improves performance, underscoring that a sparse memory set may not adequately represent the graph’s complex patterns. Nonetheless, further increasing the count of memory items beyond a certain point does not always yield better performance; it may in fact impair detection capabilities. For instance, elevating $P_{s}$ from 6 to 10 while keeping $P_{t}$ at 2 in Brain illustrates this trend, suggesting that excessively large memory modules might overemphasize specific details over general normality patterns, enabling abnormality to be reconstructed accurately.
@@ V-H2 Number of memory items
(b) Reddit
(c) Brain
Figure 7: The sensitivity of spatial memory item number $P_{s}$ and temporal memory item number $P_{t}$ on three datasets. The vertical axis represents the AUC values of STRIPE with different $P_{s}$ and $P_{t}$. A darker color indicates a higher AUC value.
@@
(b) Reddit
(c) Brain
Figure 8: The sensitivity of temporal convolution size $K_{t}$ and time window size $\tau$ on three datasets. The vertical axis represents the AUC values of STRIPE with different $K_{t}$ and $\tau$. A darker color indicates a higher AUC value.
V-H3 Parameters of temporal convolution
V-H4 Reconstruction Weight
V-H5 Hidden Dimension
In this experiment, we explored how varying the hidden dimension $D^{\prime}$ of both the encoder and decoder affects performance by adjusting $D^{\prime}$ from 8 to 256 and observing the AUC values. As shown in 9(b), we found that increasing $D^{\prime}$ enhances model performance within the range of [8, 128] for the DBLP-3, DBLP-5, and Reddit datasets, and within [8, 32] for Brain. Beyond these ranges, performance gains were negligible or even negative. Consequently, we opted for a $D^{\prime}$ of 32 for Brain and 128 for the other datasets.
V-H6 Evaluation Rounds
@@
V-I Case Study
In this section, we demonstrate the effectiveness of our spatial and temporal memory modules by visualizing the compactness loss between memory items with both normal and abnormal nodes. We configured the spatial and temporal memory items to six each and chose three normal and three abnormal nodes from the DBLP-3 test set for analysis. The compactness loss is calculated as outlined in Eq. 19. The findings are presented in Fig. 10, with darker colors indicating higher compactness loss and vice versa.
From Figs. 10(a) and 10(b), we note two key observations: (1) Normal nodes exhibit proximity to only a subset of memory items, showing distance from the rest. For instance, Fig. 10(a) illustrates that the compactness loss between the first normal node and the third spatial memory item is merely 0.06, considerably less than its loss with other memory items. This underscores the memory items’ capacity to distinctively and effectively encapsulate normal patterns, even with a limited set of items, while ensuring sufficient separation among them. (2) Conversely, abnormal nodes display substantial compactness losses across all memory items in both spatial and temporal dimensions. As an example, the compactness losses between the second abnormal node and all memory items exceed 0.89, highlighting the anomaly nodes’ divergence from established normal patterns. This deviation contributes to increased reconstruction errors and, consequently, elevated anomaly scores, effectively signaling their anomalous nature.